### v1

@@ Abstract
Understanding the interaction between humans and objects has gained much attention in recent years. Existing human-object interaction (HOI) detection methods mainly focus on the third-person perspectives, overlooking a more intuitive way from the egocentric view of HOI, namely Ego-HOI. This paper introduces an Ego-HOIBench, a new dataset to promote the benchmarking and development of Ego-HOI detection. Our Ego-HOIBench comprises more than 27K egocentric images with high-quality hand-verb-object triplet annotations across 123 fine-grained interaction categories and locations, covering a rich diversity of scenarios, object types, and hand configurations in daily activities. In addition, we explore and adapt third-person HOI detection methods to Ego-HOIBench and illustrate the challenges of hand-occluded objects and the complexity of single- and two-hand interactions. To build a new baseline, we propose a Hand Geometry and Interactivity Refinement (HGIR) scheme, which leverages hand pose and geometric information as valuable cues for interpreting interactions. Specifically, the HGIR scheme explicitly extracts global hand geometric features from the estimated hand pose proposals and refines the interaction-specific features using pose-interaction attention. This scheme enables the model to obtain a robust and powerful interaction representation, significantly improving the Ego-HOI detection capability. Our approach is lightweight and effective, and it can be easily applied to HOI baselines in a plug-and-play manner to achieve state-of-the-art results on Ego-HOIBench. Our project is available at: https://dengkunyuan.github.io/EgoHOIBench/
@@ 1 Introduction
Egocentric human-object interaction (Ego-HOI) detection aims to locate interacting human-object pairs and reason about their interaction relationships from the first-person vision. As a crucial task in human-centered advanced scene understanding, its precise outcomes can drive advancements in a wide range of downstream applications, including embodied intelligence [1, 2, 3], mixed reality [4, 5], surveillance event detection [6, 7], and visual question answering [8, 9, 10]. By analyzing images captured from a first-person perspective, the Ego-HOI assistance system offers guidance and feedback based on the user’s actions, facilitating tasks such as cooking and assembly. Furthermore, this technology has the potential to enhance embodied intelligence in imitation learning and the execution of complex tasks.
HOI detection has made significant progress in third-person vision [11, 12, 13, 14, 15, 16, 17, 18, 19]. However, it is rarely explored from the egocentric view. The primary reason is the lack of benchmark datasets clearly labeled for the Ego-HOI detection task. On the one hand, the significant domain mismatch between the HOI datasets captured from third-person vision and the egocentric task renders them unsuitable for direct application to Ego-HOI detection. As shown in Fig. 1, the third-person perspective (top row) provides a comprehensive view of the human body posture and surroundings, while the egocentric perspective (bottom row) captures interaction details of hands and objects in close range. On the other hand, although a large number of egocentric datasets have emerged in recent years, e.g., Ego4D [20], EPIC-KITCHENS[21, 22], they usually focus on the action recognition task and lack high-quality fine-grained annotations of three fundamental elements of Ego-HOI: <human hand, verb, object>. Furthermore, these datasets either only cover a single scene [22, 23, 24, 25, 26, 27] or single-hand interactions [28], or focus on rigid objects with relatively simple interaction patterns but ignore articulated objects [24, 25, 26, 28, 29], which is far from sufficient for building a comprehensive understanding of egocentric human-object interactions in real-world scenarios. The limitations of existing egocentric public datasets regarding annotation modalities and interaction diversity severely hinder the development of Ego-HOI detection.
Figure 1: Examples of human-object interactions from third-person perspective (top row) and egocentric perspective (bottom row). Different colors represent distinct elements of each HOI triplet <human/hand, verb, object>.
The narrow field of view in egocentric vision leads to severe visual occlusion [29, 30, 31], presenting a significant challenge for interaction recognition. Existing HOI detection methods are usually designed for third-person vision and rely on the rich contextual information provided by the broad view of external cameras. When applied to egocentric vision, these methods suffer from the loss of information due to mutual occlusions of hands and objects, affecting their performance in Ego-HOI detection. Due to the structural connectivity properties of human skeletons, human pose features exhibit higher robustness and reliability than traditional visual features when dealing with partial occlusions [32, 33, 34]. Based on this insight, previous studies [11, 34, 35, 36, 37, 38, 39] have attempted to incorporate pose information to distinguish subtle interaction differences. However, these methods usually depend on human pose estimators or body part detectors, which is unsuitable for hand posture estimation in egocentric scenes. Moreover, they primarily focus on extracting geometric features from the overall structure of the human body, which are not specifically designed for hands. Therefore, it is crucial to further explore flexible and effective ways to capture gesture cues to facilitate egocentric interactivity learning even under occlusion.
@@
In view of the above issues, 1) we present a new benchmark dataset, Ego-HOIBench, featuring explicit and high-quality <human hand, verb, object> triplet annotations to facilitate research on Ego-HOI perception. Our dataset covers 27,575 images and 123 hand-verb-object triplet categories, thoroughly annotating the bounding boxes and categories of human hands, active objects, and their relations. It not only extensively covers diverse hand-object interaction scenarios but also includes and distinguishes single-hand and two-hand manipulated interactions. We also define two Ego-HOIBench challenges under instance-level and image-level settings to explore the Ego-HOI detection task. 2) We propose a lightweight and effective interaction enhancement scheme, Hand Geometry and Interactivity Refinement (HGIR), that utilizes hand pose and geometric cues to improve the interaction representations from a global perspective. In particular, our approach first estimates multiple sets of candidate hand joint positions based on hand features from an HOI baseline detector. Then, we construct global hand geometric features by designing a selection strategy to identify the most suitable hand pose proposals. Next, we enhance interaction representation by pose prompts with pose-interaction attention, generating pose-aware interaction features. Finally, hand geometric features and pose-aware interaction features are fused for interaction recognition. Note that our method can be flexibly integrated with off-the-shelf HOI detectors, eliminating the need for additional hand pose estimators and achieving impressive efficiency. The main contributions of our work can be summarized as follows:
@@
  • We introduce an Ego-HOIBench, the first Ego-HOI detection benchmark containing 27K real-world egocentric images and 123 fine-grained hand-verb-object triplet annotations. Besides, we adapt and reimplement four representative third-person HOI detection methods on Ego-HOIBench, aiming to significantly advance the benchmarking works in egocentric interactive localization and recognition research.
@@
  • We propose a plug-and-play interaction enhancement scheme, i.e. HGIR, incorporating global hand pose understanding to complement and enhance interaction representations in the egocentric vision. Our approach is lightweight, effective, and general and works seamlessly with off-the-shelf HOI detection methods.
  • Experiments in representative and influential HOI baselines with our scheme validate its significant performance improvements. Extensive experiments, ablation studies, and discussions are conducted to illustrate the significance of benchmarking EgoHOI.
@@ 2.1 Egocentric Datasets and Benchmarks
With the development of wearable devices and smart glasses, increasing egocentric datasets have been proposed to study human activities from the unique first-person viewpoint [8, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 40, 41, 42]. Here, we focus on datasets related to interaction perception and localization. EPIC-KITCHENS and its extensions [21, 22] are a series of large-scale datasets that capture long-term unscripted activities in kitchen environments and densely label actions through an automatic annotation pipeline. To comprehensively study attention and action, EGTEA Gaze+ [23] records video and audio of meal preparation tasks and simultaneously provides gaze tracking, hand masks, and fine-grained action annotations. Compared with any other datasets, Ego4D [20] has a massive scale and unprecedented diversity. It builds a vast suite of benchmarks, including episodic memory, hands and objects, audio-visual diarization, social interaction, and forecasting. Datasets and benchmarks such as these fill the gaps in the first-person vision for different visual perception tasks, such as action recognition, video captioning, and hand detection.
However, most works focus on action recognition or custom tasks and do not provide joint annotations of <human hand, verb, object> required for training and evaluating Ego-HOI detection models. Many datasets [22, 30] detail all objects interacting with the hand over time but do not explicitly specify the current active object, which is essential for constructing Ego-HOI annotations. Although the MECCANO dataset [25] provides clear and complete Ego-HOI annotations, it only considers toy assembly activities. Such a limited scenario is very detrimental to the generalization of the Ego-HOI detection model. Furthermore, object diversity is essential in fully understanding interactions, but some datasets [24, 25, 26, 28, 29] tend to collect behavioral data related to rigid objects with simple interactions, ignoring articulated objects. The diversity of hand configurations is also often overlooked. The FPHA dataset [28] only presents the 3D joint locations for right hands, while H2O [29] focuses on two-hand operations. Although some datasets [23, 28] cover both single-hand and two-hand interactions, they do not consider differentiating between the left and right hands. These oversights hinder a comprehensive understanding of interactions under different hand configurations. In addition, although recent egocentric datasets have primarily focused on videos, image-based Ego-HOI detection remains highly novel and valuable for research. It is particularly well-suited for resource-limited devices and real-time applications due to its easy accessibility, fast response, and low computational requirements. The comparison between our work and existing public datasets is shown in Table 1. To the best of our knowledge, our Ego-HOIBench is the first real image-based dataset that comprises explicit and high-quality annotations of <human hand, verb, object> for Ego-HOI detection, covering a rich set of scenarios, objects, and hand configurations.
@@
Table 1: Comparison of Ego-HOIBench with existing egocentric datasets. Active Object Distinction denotes that the annotation specifies the object involved in the current interaction. Hand Dist. means distinguishing between left and right hands. (*): It can be converted from masks. (**): Only available for a subset of frames.
@@
EGTEA Gaze+ [23] Kitchens RGB 106 53 19 $\times$ $\times$ * $\times$ $\times$ 2018
EPIC-KITCHENS-100 [22] Kitchens RGB 4,053 300 97 $\times$ $\times$ 2020
Assembly101 [24] Toy assembly RGB 1,380 90 24 $\times$ $\times$ $\times$ $\times$ N/A 2022
MECCANO [25] Toy assembly RGB-D 61 20 12 $\times$ $\times$ 2022
AssemblyHands [26] Toy assembly RGB N/A N/A 6 $\times$ $\times$ $\times$ $\times$ 2023
EgoISM-HOI [27] Industrial-like RGB-D $\times$2 N/A 19 2 $\times$ 2024
@@
FPHA [28] Daily activities RGB-D 45 26 N/A $\times$ N/A $\times$ $\times$ $\times$ 2018
H2O [29] Daily activities RGB-D 36 8 11 N/A $\times$ $\times$ 2021
Ego4D [20] Multi-domain RGB N/A 871 741 N/A N/A ** $\times$ 2022
HOI4D [30] Daily activities RGB-D N/A 16 22 * $\times$ * 2022
Ego-HOIBench (Ours) Daily activities RGB-D 123 22 18 2024
@@ 2.2 HOI Detection
In recent years, HOI detection has attracted widespread research interest. This task aims to gain a fine-grained understanding of human activities by localizing human-object pairs and inferring their high-level semantic relationships. Existing HOI detection work can be categorized into two- and one-stage methods based on their detection strategies. The two-stage methods [11, 13, 15, 35, 36, 38, 43, 44, 45, 46, 47, 48] involve using a frozen object detector (e.g., Faster-RCNN [49], DETR [50]) to generate proposals for human-object pairs. These proposals are then classified using a separate network based on the features of cropped regions. The two-stage methods usually extract additional visual and contextual information, such as spatial [43, 45, 48], language [13, 15], and human pose features [11, 35, 36, 38], to improve interaction representations. Some studies [43, 47] also utilize graph structures for message propagation between detected human and object instances, thereby enhancing the reasoning performance of interactions between these instance nodes. Decoupling the stages enables training solely the interaction recognition network, thereby saving computational resources and improving training efficiency. However, optimizing the two sub-problems separately may result in suboptimal results. In contrast, the one-stage methods [16, 17, 51, 52, 53, 54, 55] directly detect HOI triplets from the entire image. Early CNN-based one-stage methods use interaction points [51] or union boxes [53] to predict interactions, but these heuristics rely on complex post-processing techniques such as Non-Maximum Suppression. Building on the success of DETR [50] in object detection, many approaches have extended the Transformer architecture to achieve end-to-end HOI detection [16, 17, 52, 54]. According to the degree of decoupling of human detection, object recognition, and interaction recognition, these methods can be further divided into single-branch [17, 56], two-branch [52], and three-branch [54] methods. Overall, these methods benefit from the strengths of Transformers in efficiently capturing long-range dependencies and have achieved significant performance improvements.

### v2

@@ Abstract
Egocentric human-object interaction (Ego-HOI) detection is crucial for intelligent agents to understand and assist human activities from a first-person perspective. However, progress has been hindered by the lack of benchmarks and methods tailored to egocentric challenges such as severe hand-object occlusion. In this paper, we introduce the real-world Ego-HOI detection task and the accompanying Ego-HOIBench, a new dataset with over 27K egocentric images and explicit, fine-grained hand-verb-object triplet annotations across 123 categories. Ego-HOIBench covers diverse daily scenarios, object types, and both single- and two-hand interactions, offering a comprehensive testbed for Ego-HOI research. Benchmarking existing third-person HOI detectors on Ego-HOIBench reveals significant performance gaps, highlighting the need for egocentric-specific solutions. To this end, we propose Hand Geometry and Interactivity Refinement (HGIR), a lightweight, plug-and-play scheme that leverages hand pose and geometric cues to enhance interaction representations. Specifically, HGIR explicitly extracts global hand geometric features from the estimated hand pose proposals, and further refines interaction features through pose-interaction attention, enabling the model to focus on subtle hand-object relationship differences even under severe occlusion. HGIR significantly improves Ego-HOI detection performance across multiple baselines, achieving new state-of-the-art results on Ego-HOIBench. Our dataset and method establish a solid foundation for future research in egocentric vision and human-object interaction understanding. Project page: https://dengkunyuan.github.io/EgoHOIBench/
@@ 1 Introduction
Egocentric human-object interaction (Ego-HOI) detection aims to accurately localize interacting human-object pairs and infer their interaction relationships from the first-person perspective. This task is fundamental for advanced human-centered scene understanding (Qiao et al., 2025) and has the potential to drive significant progress in a wide range of applications, such as embodied intelligence (Kotar & Mottaghi, 2022; Nagarajan et al., 2019; Nagarajan & Grauman, 2021), mixed reality (Wang et al., 2023; Salvato et al., 2022), surveillance event detection (Lee et al., 2024; Masuda et al., 2020), and visual question answering (Di & Xie, 2024; Fan, 2019; Jia et al., 2022). By leveraging egocentric images, Ego-HOI detection systems can provide context-aware guidance and feedback tailored to the user’s actions, enabling more intelligent assistance in daily tasks such as cooking, assembly, and navigation. Moreover, robust Ego-HOI detection is essential for advancing embodied AI, as it enables agents to better understand and imitate human behaviors, learn from demonstrations, and execute complex tasks in dynamic real-world environments.
Figure 1: Comparison between (a) third-person and (b) egocentric views of human-object interactions. The third-person view provides full-body information and a broader scene context, while the egocentric view offers a close-up, hand-centric imagery with limited context, stronger occlusions, and finer object detail—factors that necessitate different modeling strategies for Ego-HOI. Colored boxes represent different elements of each HOI triplet: human/hand, verb, and object.
Despite significant progress in HOI detection for third-person vision (Kim et al., 2022; Yuan et al., 2022; Hou et al., 2022; Park et al., 2023; Yuan et al., 2023; Zheng et al., 2023; Ning et al., 2023; Lei et al., 2023; Chen & Yanai, 2023), the egocentric perspective remains underexplored. The main barrier is the absence of benchmark datasets with explicit and high-quality annotations for Ego-HOI detection. Existing third-person HOI datasets are unsuitable for egocentric tasks due to substantial domain gaps: third-person images (Fig. 1a) provide holistic body and scene context, while egocentric images (Fig. 1b) focus on close-up hand-object interactions with limited context and frequent occlusions. Although several egocentric datasets have been introduced, such as Ego4D (Grauman et al., 2022) and EPIC-KITCHENS (Damen et al., 2018, 2022), they primarily target action recognition and lack clear joint annotation of Ego-HOI triplets, i.e., simultaneous connections between human hands, verbs, and objects within each interaction instance. Moreover, most existing datasets suffer from one or more of the following limitations: (1) coverage of only a single scene or activity domain (Li et al., 2018; Damen et al., 2022; Sener et al., 2022; Ragusa et al., 2023; Ohkawa et al., 2023; Leonardi et al., 2024); (2) focus on single-hand interactions (Garcia-Hernando et al., 2018) without distinguishing between left and right hands or supporting two-hand interactions; (3) emphasis on rigid objects, neglecting more complex articulated objects(Garcia-Hernando et al., 2018; Kwon et al., 2021; Sener et al., 2022; Ragusa et al., 2023; Ohkawa et al., 2023). These shortcomings in annotation granularity, interaction diversity, and scenario coverage severely limit the development and evaluation of robust Ego-HOI detection models.
The inherently narrow field of view in egocentric vision significantly increases the likelihood and severity of visual occlusions, as hands and manipulated objects frequently block each other or the scene (Su et al., 2025; Tekin et al., 2019; Kwon et al., 2021; Liu et al., 2022). This presents a major challenge for interaction recognition, as critical cues for understanding human-object interactions are often partially or fully hidden. Most existing HOI detection methods are developed for third-person vision, where a wide field of view provides abundant contextual and spatial information, reducing the impact of occlusion. When these methods are applied to egocentric settings, their performance degrades due to the loss of contextual cues and the prevalence of mutual occlusions between hands and objects. As a result, Ego-HOI detection requires more robust feature representations that can handle missing or ambiguous visual information. Human pose features, particularly those derived from geometry, have demonstrated greater robustness to partial occlusion due to the structural connectivity of the human skeleton (Qiao et al., 2022; Tu et al., 2023; Rajasegaran et al., 2023). Prior works (Gupta et al., 2019; Wan et al., 2019; Zheng et al., 2020; Qiao et al., 2022; Li et al., 2022; Park et al., 2023) have explored incorporating pose information to better distinguish subtle interaction differences. However, these approaches typically rely on external pose estimators trained on third-person data, which are not optimized for egocentric viewpoints and can hardly capture the fine-grained hand cues critical for Ego-HOI understanding. This highlights the urgent need for methods that can flexibly and accurately capture hand pose information directly within the egocentric context, enabling more robust and reliable interaction recognition.
@@
In view of the above issues, 1) we introduce Ego-HOIBench, a comprehensive and rigorously annotated benchmark specifically tailored for Ego-HOI detection. Ego-HOIBench provides explicit <human hand, verb, object> triplet annotations for 27,575 real-world images, covering 123 diverse hand-verb-object triplet categories. Each image is thoroughly annotated with bounding boxes and categories for both hands and active objects, as well as their interaction relationships. Our dataset uniquely captures a wide range of egocentric scenarios and objects, including both single-hand and two-hand manipulated interactions, offering an unprecedented testbed for advancing Ego-HOI research. 2) We propose Hand Geometry and Interactivity Refinement (HGIR), a novel, lightweight, and plug-and-play scheme that leverages global hand pose and geometric cues to significantly enhance interaction representations. Our approach first estimates multiple candidate hand joint positions from hand features produced by any HOI baseline detector, eliminating the need for external pose estimators. We then construct robust global hand geometric features through a designed selection strategy, and refine interaction representations using pose-interaction attention to generate pose-aware features. Finally, we fuse hand geometric and pose-aware interaction features to achieve superior interaction recognition. Our method is highly efficient, generalizable, and can be seamlessly integrated with existing HOI detectors. The main contributions of our work are summarized as follows:
@@
  • We introduce the Ego-HOI Detection task and Ego-HOIBench dataset, the first real-world Ego-HOI detection benchmark with explicit joint hand-verb-object triplet annotations, providing a comprehensive and realistic testbed for advancing Ego-HOI research. Besides, we adapt and reimplement four representative third-person HOI detection methods on Ego-HOIBench, aiming to significantly advance the benchmarking works in egocentric interactive recognition and detection research.
@@
  • We propose a plug-and-play interaction enhancement scheme, i.e. HGIR, incorporating global hand pose understanding to complement and enhance interaction representations in the egocentric vision. Our approach is lightweight, effective and works seamlessly with off-the-shelf HOI detection methods, enabling substantial performance gains with minimal computational overhead.
  • We conduct extensive experiments by integrating our scheme with representative HOI detection baselines, demonstrating consistent and significant performance improvements. Comprehensive ablation studies and analyses further highlight the importance of benchmarking Ego-HOI detection and the effectiveness of our approach.
@@ 2.1 Egocentric Datasets and Benchmarks
With the rise of wearable devices and smart glasses, such as Apple Vision Pro (Wienrich & Obremski, 2025) and Microsoft Hololens 2 (Balakrishnan & Guo, 2024), the field of egocentric vision has witnessed a surge in dedicated datasets aimed at understanding human activities from the first-person perspective (Damen et al., 2018; Li et al., 2018; Garcia-Hernando et al., 2018; Kwon et al., 2021; Jia et al., 2022; Grauman et al., 2022; Liu et al., 2022; Damen et al., 2022; Zhang et al., 2022b, c; Sener et al., 2022; Ragusa et al., 2023; Ohkawa et al., 2023; Leonardi et al., 2024). Among these, datasets focusing on interaction perception and localization have become particularly important for advancing egocentric understanding. For example, the EPIC-KITCHENS series (Damen et al., 2018, 2022) provides large-scale, densely annotated recordings of unscripted kitchen activities, enabling comprehensive study in realistic environments. EGTEA Gaze+ (Li et al., 2018) enriches egocentric research by offering video, audio, gaze tracking, and detailed action annotations, supporting multi-modal analysis of attention and interaction. Ego4D (Grauman et al., 2022) stands out for its unprecedented scale and diversity, introducing a broad suite of benchmarks that cover episodic memory, audio-visual diarization, and social interaction.
Despite the progress in egocentric vision datasets, there remain significant limitations in providing the joint annotations necessary for Ego-HOI detection models. Most existing datasets focus on action recognition or custom tasks and lack explicit, simultaneous annotations of <human hand, verb, object> triplets required for comprehensive Ego-HOI detection (Grauman et al., 2022). For example, many datasets (Damen et al., 2022; Liu et al., 2022) annotate all objects interacting with the hand over time but do not specify which object is actively involved in the current interaction, making it difficult to construct precise Ego-HOI labels. The MECCANO dataset (Ragusa et al., 2023) is one of the few that provides clear active object annotations, but it is limited to toy assembly activities, restricting the generalizability of models trained on it. Furthermore, object and interaction diversity are often lacking: several datasets (Garcia-Hernando et al., 2018; Kwon et al., 2021; Sener et al., 2022; Ragusa et al., 2023; Ohkawa et al., 2023) focus mainly on rigid objects and simple interactions, neglecting articulated objects with more complex interactions. Hand configuration diversity is also insufficiently addressed; for instance, FPHA (Garcia-Hernando et al., 2018) only provides 3D joint locations for right hands, while H2O (Kwon et al., 2021) emphasizes two-hand operations without distinguishing between left and right hands. Even datasets that include both single- and two-hand interactions (Li et al., 2018; Garcia-Hernando et al., 2018) often fail to differentiate hand sides, limiting their utility for detailed Ego-HOI modeling. These gaps in annotation granularity and diversity hinder the development and evaluation of robust Ego-HOI detection models capable of handling real-world complexity.
Table 1 summarizes the comparison between our work and existing public datasets. To the best of our knowledge, Ego-HOIBench is the first real image-based dataset to provide explicit and high-quality joint annotations of <human hand, verb, object> for Ego-HOI detection, covering a rich diversity of scenarios, objects, and hand configurations, and directly addressing the limitations of prior datasets.
@@
Table 1: Comparison of Ego-HOIBench and existing egocentric datasets in terms of scenario diversity, annotation modalities, and interaction details. “Active Object Distinction” indicates whether the dataset explicitly annotates the object currently involved in the interaction. “Hand Dist.” denotes whether left and right hands are distinguished. (*) means can be derived from masks. (**) means available only for a subset of frames.
@@
EGTEA Gaze+ (Li et al., 2018) Kitchens RGB 106 53 19 $\times$ $\times$ * $\times$ $\times$ 2018
EPIC-KITCHENS-100 (Damen et al., 2022) Kitchens RGB 4,053 300 97 $\times$ $\times$ 2020
Assembly101 (Sener et al., 2022) Toy assembly RGB 1,380 90 24 $\times$ $\times$ $\times$ $\times$ N/A 2022
MECCANO (Ragusa et al., 2023) Toy assembly RGB-D 61 20 12 $\times$ $\times$ 2022
AssemblyHands (Ohkawa et al., 2023) Toy assembly RGB N/A N/A 6 $\times$ $\times$ $\times$ $\times$ 2023
EgoISM-HOI (Leonardi et al., 2024) Industrial-like RGB-D $\times$2 N/A 19 2 $\times$ 2024
@@
FPHA (Garcia-Hernando et al., 2018) Daily activities RGB-D 45 26 N/A $\times$ N/A $\times$ $\times$ $\times$ 2018
H2O (Kwon et al., 2021) Daily activities RGB-D 36 8 11 N/A $\times$ $\times$ 2021
Ego4D (Grauman et al., 2022) Multi-domain RGB N/A 871 741 N/A N/A ** $\times$ 2022
HOI4D (Liu et al., 2022) Daily activities RGB-D N/A 16 22 * $\times$ * 2022
Ego-HOIBench (Ours) Daily activities RGB-D 123 22 18 2025
@@ 2.2 HOI Detection
In recent years, HOI detection has attracted widespread research interest. This task aims to gain a fine-grained understanding of human activities by both localizing human-object pairs and inferring their high-level semantic relationships. Existing HOI detection work are generally divided into two- and one-stage methods based on their detection strategies. The two-stage methods (Wan et al., 2019; Gao et al., 2020; Zheng et al., 2020; Z. Zhang et al., 2021; Yang et al., 2021; Li et al., 2022; Zhang et al., 2022a, d; Park et al., 2023; Zheng et al., 2023; Lei et al., 2023; Zhang et al., 2023) first employ a dedicated object detector (e.g., Faster-RCNN (Ren et al., 2016), DETR (Carion et al., 2020)) to generate candidate human-object pairs, and then classify the interactions between these pairs using a separate recognition network. This decoupled pipeline allows for the integration of rich contextual cues, such as spatial relationships (Yang et al., 2021; Gao et al., 2020; Zhang et al., 2022a), language priors (Zheng et al., 2023; Lei et al., 2023), and human pose features (Wan et al., 2019; Zheng et al., 2020; Li et al., 2022; Park et al., 2023), to enhance interaction understanding. Furthermore, several works (Gao et al., 2020; Z. Zhang et al., 2021; Hong et al., 2025) utilize graph structures for message propagation between detected human and object instances, thereby enhancing the reasoning performance of interactions between these instance nodes. Decoupling the stages enables training solely the interaction recognition network, thereby saving computational resources and improving training efficiency. However, optimizing the two sub-problems separately may result in suboptimal results, and the reliance on the object detectors can introduce error propagation.
In contrast, one-stage methods (Kim et al., 2020, 2021, 2022; Chen & Yanai, 2023; Lin et al., 2023; Kim et al., 2023) aim to directly predict HOI triplets from the entire image in a unified framework. Early CNN-based one-stage approaches, such as those using interaction points (Liao et al., 2020) or union boxes (Kim et al., 2020), often depend on heuristic designs and require complex post-processing steps like Non-Maximum Suppression, which may limit their scalability and robustness. The advent of DETR (Carion et al., 2020) and its Transformer-based architecture has inspired a new generation of end-to-end HOI detectors (Chen & Yanai, 2023; Kim et al., 2022, 2021, 2023), which eliminate the need for hand-crafted post-processing and enable joint optimization of detection and interaction recognition. These methods can be further categorized by their architectural design: single-branch (Tamura et al., 2021; Kim et al., 2022), two-branch (Kim et al., 2021), and three-branch (Kim et al., 2023), depending on how they decouple human, object, and interaction prediction. One-stage approaches offer significant advantages in inference efficiency and end-to-end learning, and have demonstrated strong performance on large-scale benchmarks. However, they may require more training data to generalize well and are less flexible in integrating external cues compared to two-stage pipelines.
The choice between one-stage and two-stage HOI detection methods depends on application needs. To benchmark Ego-HOIBench, we adapt and reimplement several representative and influential third-person HOI detectors in the egocentric setting, including two-stage method STIP (Zhang et al., 2022d), one-stage one-branch method QPIC (Tamura et al., 2021), one-stage two-branch method HOTR (Kim et al., 2021), and one-stage three-branch method MUREN (Kim et al., 2023)).